# Stabilizing Non-Abelian Topological Order against Heralded Noise via Local Lindbladian Dynamics

### v1

@@ III.2 Correction dynamics for the Toric code
We consider a honeycomb lattice on the 2d torus, with a qubit (together with an $X$ and a $Z$ flag) placed on each edge. The periodic boundary conditions (PBC) on the $L_{x}\times L_{y}$ sized torus are implemented by making the following identifications for the position vector $\vec{r}$:
for $n_{x},n_{y}\in\mathbb{Z}$ and where $\vec{a},\vec{b}$, and $\vec{c}$ are the basis vectors of the honeycomb lattice, as shown in Fig. 2; these are related by $\vec{a}=\vec{b}+\vec{c}$. In the following, we set $L_{x}=L_{y}=L$ which results in lattice with $L^{2}$ edges, $L^{2}/3$ plaquettes, and $32L^{2}/$ 3 vertices. Note that the Toric code can be defined on any 2D lattice; we choose the honeycomb lattice here to make contact with the $D_{4}$ model discussed later.
The Toric code ground states |$\ket{\psi}$ are those states that are simultaneously stabilized by a set of vertex $B_{v}$ and plaquette $A_{p}$ stabilizers i.e., they satisfy $B_{v}\ket{\psi}=A_{p}\ket{\psi}=\ket{\psi}\,\forall\,v,p$, with
$B_{v}=\prod_{j\in^{*}v}Z_{j}\,,\qquad A_{p}=\prod_{k\in\partial p}X_{k}\,.$ (6)
A key feature of the stabilizers in Eq. (6) is that vertex (plaquette) stabilizer defects are necessarily created in pairs, and are spatially separated via a string of contiguous Pauli-$X$ (Pauli-$Z$) errors. This is a robust feature of the model, because the stabilizer defects correspond to emergent anyons, which cannot be locally created. For perfectly heralded errors, this means that pairs of stabilizer defects are necessarily connected by strings of flags. Our correction protocol uses the information provided by the positions of these flags to appropriately move stabilizer defects along such strings, thereby annihilating each stabilizer defect with its partner.
Figure 3: $X$ correction protocol for Toric code: The edges that have $X$ flags on them are shown using bold lines and the vertex defects are depicted by black triangles. The initial configurations on the left side are updated to obtain the configuration on the right side. We illustrate processes in which stabilizer defects are present near the corrected vertex; the processes with no stabilizer defects act identically on the flags, but do not act on the qubits. (a) X-leaf move: Vertex $v$ has a single edge with an $X$ flag on it, and hosts a stabilizer defect. The defect is displaced by applying $X_{j}$ on the flagged edge $j$, after which the $X$ flag is removed. (b) X-loop move: (i) The two edges ($e_{0}$ and $e_{1}$) emanating from a vertex $v$ of the purple sublattice and bordering its northwest plaquette $p_{NW}$ have $X$ flags on them. The correction step in Eq. (8) removes these flags, raises new $X$ flags on the remaining edges of $p_{NW}$ (if they were not present to begin with; highlighted in green color) and displaces the vertex defect by applying $X_{e_{0}}$. (ii) The X-loop move can also be applied at vertex $\widetilde{v}$ which locally hosts a flag configuration $n^{f}_{X_{e_{0}}}=n^{f}_{X_{e_{1}}}=1$, and $n^{f}_{X_{e_{2}}}=0$ even though the flags do not form a closed loop. As a result of this operation, the vertex defects will be paired along the alternative path highlighted in green color. The loop moves are applied only at vertices on the purple sublattice.
@@
To see how the correction protocol works in practice, let us consider a vertex stabilizer defect $B_{v}=-1$. If only one of the three edges (say edge $j$) surrounding this defect has an $X$ flag, then we can be certain that the defect is result of an $X$ type error that occurred at this particular qubit. Consequently, we move the defect by applying an $X$ operator to this qubit and simultaneously lower the $X$ flag, as illustrated in Fig. 3(a). Similarly, at a vertex with no stabilizer defect and where only one of the edges has an $X$ flag, we can simply erase the flag. These moves, which we call leaf moves, are implemented via the local jump operators:
@@
Here, $v$ labels the $32L^{2}/$ 3 vertices of the honeycomb lattice and $k\in^{*}v\setminus j$ denotes the two edges $\neq j$ attached to vertex $v$. The scalar $s=-1(s=+1)$ corresponds to the case where a vertex defect is present (absent) and $\gamma_{x}$ parameterizes the local correction rate. The dynamics generated by this jump operator autonomously implements measurement and feedback towards a defect free state (and requires no postselection).
@@
Consider the configuration of flags and defects shown in the left panel of Fig. 3(b). Here, the vertex $v$ on the purple sublattice is found with flags on the two edges $e_{0}$ and $e_{1}$ that are North or West of it, and no flag on the third edge $e_{2}$. Our protocol detects such configurations by checking if the flag configuration is of the form $n^{f}_{X_{e_{0}}}=n^{f}_{X_{e_{1}}}=1$ and $n^{f}_{X_{e_{2}}}=0$. When such a configuration is detected, the protocol lowers the flags on $e_{0}$ and $e_{1}$, and raises new $X$ flags on the outer edges of the North-Western plaquette $p_{NW}$ (if they are not already present, as is the case for the vertex $\widetilde{v}$ in the Figure). The $B_{v}$ stabilizer at the central vertex is then measured, and if $B_{v}=-1$, the defect is moved up and to the right by applying $X_{e_{0}}$. The overall result of this operation is to push the string or loop of $n^{f}_{X}=1$ flags in the North-West direction. Concurrently, the vertex defects are (a) either paired-up along the path or (b) the newly raised flags provide an alternate path on the Honeycomb lattice that is equivalent to the original error-string up to a product of plaquette stabilizers $A_{p}$. Since the purpose of these moves is to shrink the loop in a preferential direction, the $X$-loop moves are not applied to vertices on yellow sublattice. At a vertex $v$ on the purple sublattice, this operation is implemented by the local jump operators:
@@
Figure 4: $Z$ correction protocol for Toric code: The edges that have $Z$ flags are shown using bold wavy lines. The correction operation is applied to the blue plaquette in the top row to obtain the new configuration in the bottom row. Here, we illustrate processes in which stabilizer defects are moved by acting with $Z$ on appropriate edges (shown using dashed arrows); the processes with no stabilizer defects act identically on the flags, but do not apply the Pauli-$Z$ operator. (a) Z-leaf move: a plaquette defect occurs on a plaquette with a single flagged edge. The correction protocol applies $Z$ to this flagged edge, and subsequently erases the flag, moving the defect to a neighboring plaquette. (b) Z-loop moves: Four candidate processes corresponding to $Z$ flags on edges labeled by $(e_{0},e_{1})$,$(e_{1},e_{2})$,$(e_{0},e_{2})$, and $(e_{0},e_{1},e_{2})$ are shown here. In each configuration, new $Z$ flags are added on edges $d_{0}$ and/or $d_{1}$ as described in the main text (refer to Eq. (10)).
@@
As with $X$ flags, loop moves are also necessary to remove $Z$ flags that form closed loops on the dual lattice. Let us consider a plaquette $p$ that has more than one $Z$ flag on its North-Western edges. The possible configurations can be labeled by the occupied edges =$\mathcal{E}=\{(e_{0},e_{1}),(e_{1},e_{2}),(e_{0},e_{2}),(e_{0},e_{1},e_{2})\}$, as illustrated in Fig. 4(b). For each configuration, the figure describes the locations $R[\mathcal{E}]=\{(d_{0}),(d_{1}),(d_{0},d_{1}),(d_{0},d_{1})\}$ of new $Z$ flags which are consistent with this network, meaning that errors on these edges are equivalent to errors on the original flagged edges, up to a product of vertex stabilizers $B_{v}$. For a given $\varepsilon\in\mathcal{E}$ and plaquette $p$, the corresponding jump operator is:
@@
Here, $s=\pm 1$ denotes the stabilizer measurement outcome and $\mathbf{m^{z}}\in\{0,1\}^{\otimes\text{dim}(R[\varepsilon])}$ is bit-string that accounts for all possible configurations of $Z$ flags on the edges in $R[\varepsilon]$. Upon detection of plaquette defect, i.e. $s=-1$, it is always moved towards north-most neighbor which shares the edge $\varepsilon^{(0)}$ with $p$.
The $X$ ($Z$) loop moves allow the dynamics to shrink contractible loops of flags that may span multiple plaquettes (or vertices, for $Z$ flags). By choosing a particular direction in which to “push” loop segments, this can be accomplished using the sequence of local moves acting on a single plaquette (vertex) and the edges in its immediate vicinity. Evidently, this also means that the loop moves do not act only on loops; in certain configurations they will instead re-arrange the local flag and stabilizer defect configuration, as shown in Fig. 3(b-ii). Note however that each of these moves only involve an $O(1)$ number of flags and stabilizer operators, thereby retaining the local nature of the dynamics.
@@
$\begin{split}\mathcal{L}_{C}=&\ \gamma_{x}\sum_{\begin{subarray}{c}v,s=\pm 1\\ j\in^{*}v\end{subarray}}\mathcal{L}^{\text{TC}^{\text{X-leaf}}}_{v,j,s}\ +\ \gamma_{z}\sum_{\begin{subarray}{c}p,s=\pm 1\\ j\in\partial p\end{subarray}}\mathcal{L}^{\text{TC}^{\text{Z-leaf}}}_{p,j,s}\\ &+\gamma_{x}\sum_{\begin{subarray}{c}v\in{\text{purple}},s=\pm 1\\ \mathbf{m^{x}}\in\{0,1\}^{\otimes 4}\end{subarray}}\mathcal{L}^{\text{TC}^{\text{X-loop}}}_{v,s,\mathbf{m^{x}}}+\gamma_{z}\sum_{\begin{subarray}{c}p,s=\pm 1,\\ \varepsilon\in\mathcal{E}(p),\mathbf{m^{z}}\end{subarray}}\mathcal{L}^{\text{TC}^{\text{Z-loop}}}_{p,s,\varepsilon,\mathbf{m^{z}}}\,,\end{split}$ (11)
@@
Figure 5: Time evolution in the Toric code on a $48\times 48$ honeycomb lattice for four representative points in the $\gamma_{x}$ $\gamma_{z}$ phase diagram (see Fig. 1(a)). Plots show flag densities $n^{f}_{X,Z}$, densities of vertex $n^{d}_{B}$ and plaquette $n^{d}_{A}$ defects, and probability of logical errors as a function of time. The state of logical information is determined by pairing the stabilizer defects in the time evolved state using the MWPM decoder implemented in the PyMatching-2 package [102]. The probabilities of logical $X$ errors ($\equiv 1-F_{|\widetilde{0}\rangle_{L}}$) and logical $Z$ errors ($\equiv 1-F_{|\widetilde{+}\rangle_{L}}$) are computed in terms of the fidelity of this decoded state and two distinct initial states defined in Eq. (14). The noise rate is set to $\eta=1$, with each data point representing an average over $500$ Monte Carlo runs. (a) Absorbing phase $(\gamma_{x}=4,\gamma_{z}=8)$: when both correction rates are small, flags rapidly reach their absorbing state $n^{f}_{X}=n^{f}_{Z}=1$, and stabilizer defect densities tend to $n^{d}_{A}=n^{d}_{B}=1/2$, indicating that all qubits become maximally mixed. The logical error rates $1-F_{|\widetilde{0}\rangle_{L}}$ and $1-F_{|\widetilde{+}\rangle_{L}}$ both saturate to their maximal value and hence no information is retained in the steady-state. (b) Partially X-active phase $(\gamma_{x}=15,\gamma_{z}=8)$: The density of $X$ flags and $B$ defects saturates at a small value, resulting in a small logical $X$ error rate ( $1-F_{|\widetilde{0}\rangle_{L}}$) which protects the $Z$-logical information, whereas $Z$ flags and $A$ defects proliferate, rapidly destroying the $X$- logical information. (c) The analogous partially Z-active phase $(\gamma_{x}=4,\gamma_{z}=20)$, where only logical $X$ information is protected as a result of a small probability of logical $Z$ errors ($1-F_{|\widetilde{+}\rangle_{L}}$). (d) Active phase $(\gamma_{x}=15,\gamma_{z}=20)$: here, both $X$ and $Z$ flags saturate to small values. Both logical $Z$ (cyan colored curve) and logical $X$ (magenta colored curve) error rates are close to zero and hence the quantum information can be recovered up to times exponentially large in system size in this phase.
@@ III.3 Metrics for topological order and numerical results
The full Hilbert space of our model, consisting of both qubits and flags, is spanned by the combined eigenstates of the $A_{p},B_{v}$ stabilizers and the flag occupation-numbers. The noise and correction Lindbladians in Eq. (12), which generate the total dynamics, do not couple the diagonal and off-diagonal elements of the density matrix $\rho$ in this basis. Hence, the time evolution can be simulated as a stochastic Monte Carlo dynamics of stabilizer and flag populations [103]. We initialize the system in the Toric code ground state subspace, where $B_{v}=A_{p}=+1$ for all vertices $v$ and plaquettes $p$, with the flags set to $n^{f}_{X_{j}}=n^{f}_{Z_{j}}=0$ on all edges $j$. We then simulate the dynamics using random sequential updates of the state, where the rates of all processes are specified in the previous section. The time evolved observables are computed by averaging over multiple independent Monte Carlo runs (see Appendix. A for additional details). In the what follows, we set the noise rate of the erasure channel to $\eta=1$ and explore the steady-state phase diagram as a function of the $X$ and $Z$ correction rates $\gamma_{x}$ and $\gamma_{z}$, respectively.
@@
In the active phase, the stabilizer defects are corrected along the strings of erasures frequently enough such that the density of defects remains small, as shown in Fig. 5(d) at a representative point in the phase. Here, the densities of vertex and plaquette stabilizer defects are defined respectively as
@@
The heralded nature of the noise and the correction dynamics also ensures that pairs of vertex (plaquette) defects are necessarily connected by strings of $X$ ($Z$) type flags. Hence, a low density of flags necessarily corresponds to confinement of the stabilizer defects. We further illustrate this by evaluating the probability of a logical error upon pairing the defects in the time-evolved state using the MWPM decoder, ignoring the configuration of the flags. The logical state (i.e. the exact initial state within the ground state Hilbert space) is recovered by this decoder if the probability of both logical $X$ errors (logical bit-flips) and logical $Z$ errors (logical phase-flips) remains low. These error probabilities can be estimated by initializing the system in two different basis states in the Toric code ground-state space defined as
@@
where $\mathcal{Z}_{V/H}$ are the Wilson line operators shown in Fig. 2 (which also serve as logical Pauli-$Z$ operators), and the logical Pauli-$X$ operators $\mathcal{X}_{H/V}$ are equivalent line operators comprised of products of $X$ operators along closed vertical and horizontal curves on the direct lattice. The state of the qubits in a particular Monte Carlo trajectory $r$ initialized in $|\psi_{t=0}\rangle$ can be written as
@@
has no stabilizer defects. If the edges in $e^{(r)}_{x}\oplus d^{(r)}_{x}$ ($e^{(r)}_{z}\oplus d^{(r)}_{z}$) form an odd number of non-contractible loops around the torus, they act as a logical $X$ (logical $Z$) operator on the state. Since $\{\mathcal{Z}_{V},\mathcal{X}_{H}\}=\{\mathcal{Z}_{H},\mathcal{X}_{V}\}=0$, such non-contractible loops flip the eigenvalue of the logical $Z$ (logical $X)$ operator relative to the initial state, thereby taking $|\widetilde{0}\rangle_{L}\leftrightarrow|\widetilde{1}\rangle_{L}$ (|+~$|\widetilde{+}\rangle_{L}\leftrightarrow|\widetilde{-}\rangle_{L}$). We characterize the probability of such logical flips in terms of fidelity of decoded state with respect to the initial state as
@@
where $N_{r}$ is number of independent Monte Carlo realizations. The probability of logical $X$ error (logical $Z$ error) is then equivalent to $1-F_{|\widetilde{0}\rangle_{L}}$ ($1-F_{|\widetilde{+}\rangle_{L}}$). In Fig. 5(d), we see that the probability of both types of logical flips is essentially zero in the active state. This provides explicit evidence that quantum information is protected by the dynamical correction protocol for long times.
In contrast, the absorbing state is the unique dead state for the flags: when the flag density is $1$, none of our correction operations can act on any sites and the dynamics consequently no longer acts on the flags. In this state, since all the correction jump operators vanish identically, the qubits evolve under the depolarizing channel and the dynamics leads to a maximally mixed steady state. This is evident in Fig. 5(a), which shows that the stabilizer defect density in this phase saturates to $1/2$. All encoded information is consequently lost, as is evident from the fact that the probability of logical flips saturates to $0.751-\frac{1}{2^{2}}=$ 0.75 since the decoder returns all possible outcomes with equal probability.
In the partially $X$-active ($Z$-active) phase, only $Z$ (only $X$) flags have proliferated. Thus, our correction protocol no longer corrects $Z$ ($X$) errors, but remains effective on $X$ ($Z$) errors. One manifestation of this is that only logical $Z$ (logical $X$) information is protected in the steady-state, as shown in Fig. 5(b) and (c). Each of these phases thus represents a classical memory, in which the information encoded in only one type of logical operator is dynamically protected. These phases represent intrinsically mixed topological phases which cannot be realized in the ground-states of 2d local gapped Hamiltonians [40, 55] since they cannot be two-way channel connected to the trivial state on the 2-torus (even though are two-way channel connected on the infinite plane).

### v2

@@ III.2 Correction dynamics for the Toric code
A key feature of the Toric code stabilizers in Eq. (3) is that vertex (plaquette) stabilizer defects are necessarily created in pairs, and are spatially separated via a string of contiguous Pauli-$X$ (Pauli-$Z$) errors. This is a robust feature of the model because the stabilizer defects correspond to emergent anyons, which cannot be locally created. For perfectly heralded errors, this means that pairs of stabilizer defects are necessarily connected by strings of flags. Our correction protocol uses the information provided by the positions of these flags to appropriately move stabilizer defects along such strings, thereby annihilating each stabilizer defect with its partner.
Figure 3: $X$ correction protocol for Toric code: The edges that have $X$ flags on them are shown using bold lines, and the vertex defects are depicted by black triangles. The initial configurations on the left side are updated to obtain the configuration on the right side. We illustrate processes in which stabilizer defects are present near the corrected vertex; the processes with no stabilizer defects act identically on the flags, but do not act on the qubits. (a) X-leaf move: Vertex $v$ has a single edge with an $X$ flag on it, and hosts a stabilizer defect. The defect is displaced by applying $X_{j}$ on the flagged edge $j$, after which the $X$ flag is removed. (b) X-loop move: (i) The two edges ($e_{0}$ and $e_{1}$) emanating from a vertex $v$ of the purple sublattice and bordering its northwest plaquette $p_{NW}$ have $X$ flags on them. The correction step in Eq. (9) removes these flags, raises new $X$ flags on the remaining edges of $p_{NW}$ (if they were not present to begin with; highlighted in green color), and displaces the vertex defect by applying $X_{e_{0}}$. (ii) The X-loop move can also be applied at vertex $\widetilde{v}$ which locally hosts a flag configuration $n^{f}_{X_{e_{0}}}=n^{f}_{X_{e_{1}}}=1$, and $n^{f}_{X_{e_{2}}}=0$ even though the flags do not form a closed loop. As a result of this operation, the vertex defects will be paired along the alternative path highlighted in green color. The loop moves are applied only at the vertices on the purple sublattice.
@@
To see how the correction protocol works in practice, let us consider a vertex stabilizer defect $B_{v}=-1$. If only one of the three edges (say edge $j$) surrounding this defect has an $X$ flag, then we can be certain that the defect is the result of an $X$ type error that occurred at this particular qubit. Consequently, we move the defect by applying an $X$ operator to this qubit and simultaneously lower the $X$ flag, as illustrated in Fig. 3(a). Similarly, at a vertex with no stabilizer defect and where only one of the edges has an $X$ flag, we can simply erase the flag. These moves, which we call leaf moves, are implemented via the local jump operators:
@@
Here, $v$ labels the $2L^{2}/3$ vertices of the honeycomb lattice and $k\in^{*}v\setminus j$ denotes the two edges $\neq j$ attached to vertex $v$. The scalar s=1(s=+1)s=-1(s=+1) corresponds to the case where a vertex defect is present (absent), and $\gamma_{x}$ parametrizes the local correction rate. The dynamics generated by this jump operator autonomously implements measurement and feedback towards a defect-free state (and requires no postselection).
@@
Consider the configuration of flags and defects shown in the left panel of Fig. 3(b). Here, the vertex $v$ on the purple sublattice is found with flags on the two edges $e_{0}$ and $e_{1}$ that are North or West of it, and no flag on the third edge $e_{2}$. Our protocol detects such configurations by checking if the flag configuration is of the form $n^{f}_{X_{e_{0}}}=n^{f}_{X_{e_{1}}}=1$ and $n^{f}_{X_{e_{2}}}=0$. When such a configuration is detected, the protocol lowers the flags on $e_{0}$ and $e_{1}$, and raises new $X$ flags on the outer edges of the North-Western plaquette $p_{NW}$ (if they are not already present, as is the case for the vertex $\widetilde{v}$ in the Figure). The $B_{v}$ stabilizer at the central vertex is then measured, and if $B_{v}=-1$, the defect is moved up and to the right by applying $X_{e_{0}}$. The overall result of this operation is to push the string or loop of $n^{f}_{X}=1$ flags in the North-West direction. Concurrently, the vertex defects are (a) either paired-up along the path or (b) the newly raised flags provide an alternate path on the Honeycomb lattice that is equivalent to the original error-string up to a product of plaquette stabilizers $A_{p}$. Since the purpose of these moves is to shrink the loop in a preferential direction, the $X$-loop moves are not applied to vertices on the yellow sublattice. At a vertex $v$ on the purple sublattice, this operation is implemented by the local jump operators:
@@
Figure 4: $Z$ correction protocol for Toric code: The edges that have $Z$ flags are shown using bold wavy lines. The correction operation is applied to the blue plaquette in the top row to obtain the new configuration in the bottom row. Here, we illustrate processes in which stabilizer defects are moved by acting with $Z$ on appropriate edges (shown using dashed arrows); the processes with no stabilizer defects act identically on the flags, but do not apply the Pauli-$Z$ operator. (a) Z-leaf move: a plaquette defect occurs on a plaquette with a single flagged edge. The correction protocol applies $Z$ to this flagged edge, and subsequently erases the flag, moving the defect to a neighboring plaquette. (b) Z-loop moves: Four candidate processes corresponding to $Z$ flags on edges labeled by $(e_{0},e_{1})$$(e_{1},e_{2})$$(e_{0},e_{2})$, and $(e_{0},e_{1},e_{2})$ are shown here. In each configuration, new $Z$ flags are added on edges $d_{0}$ and/or $d_{1}$ as described in the main text (refer to Eq. (11)).
@@
As with $X$ flags, loop moves are also necessary to remove $Z$ flags that form closed loops on the dual lattice. Let us consider a plaquette $p$ that has more than one $Z$ flag on its North-Western edges. The possible configurations can be labeled by the occupied edges $\mathcal{E}=\{(e_{0},e_{1}),(e_{1},e_{2}),(e_{0},e_{2}),(e_{0},e_{1},e_{2})\}$, as illustrated in Fig. 4(b). For each configuration, the figure describes the locations $R[\mathcal{E}]=\{(d_{0}),(d_{1}),(d_{0},d_{1}),(d_{0},d_{1})\}$ of new $Z$ flags which are consistent with this network, meaning that errors on these edges are equivalent to errors on the original flagged edges, up to a product of vertex stabilizers $B_{v}$. For a given $\varepsilon\in\mathcal{E}$ and plaquette $p$, the corresponding jump operator is:
@@
Here, $s=\pm 1$ denotes the stabilizer measurement outcome and $\mathbf{m^{z}}\in\{0,1\}^{\otimes\text{dim}(R[\varepsilon])}$ is bit-string that accounts for all possible configurations of $Z$ flags on the edges in $R[\varepsilon]$. Upon detection of a plaquette defect, i.e. s=1s=-1, it is always moved towards the north-most neighbor which shares the edge $\varepsilon^{(0)}$ with $p$.
The $X$ ($Z$) loop moves allow the dynamics to shrink contractible loops of flags that may span multiple plaquettes (or vertices, for $Z$ flags). By choosing a particular direction in which to “push” loop segments, this can be accomplished using the sequence of local moves acting on a single plaquette (vertex) and the edges in its immediate vicinity. Evidently, this also means that the loop moves do not act only on loops; in certain configurations, they will instead re-arrange the local flag and stabilizer defect configuration, as shown in Fig. 3(b-ii). Note, however, that each of these moves only involves an $O(1)$ number of flags and stabilizer operators, thereby retaining the local nature of the dynamics.
@@
$\mathcal{L}_{C}$ $=\ \gamma_{x}\sum_{\begin{subarray}{c}v,s=\pm 1\\ j\in^{*}v\end{subarray}}\mathcal{L}^{\text{TC}^{\text{X-leaf}}}_{v,j,s}\ +\ \gamma_{z}\sum_{\begin{subarray}{c}p,s=\pm 1\\ j\in\partial p\end{subarray}}\mathcal{L}^{\text{TC}^{\text{Z-leaf}}}_{p,j,s}$ (12)
$+\gamma_{x}\sum_{\begin{subarray}{c}v\in{\text{purple}},s=\pm 1\\ \mathbf{m^{x}}\in\{0,1\}^{\otimes 4}\end{subarray}}\mathcal{L}^{\text{TC}^{\text{X-loop}}}_{v,s,\mathbf{m^{x}}}+\gamma_{z}\sum_{\begin{subarray}{c}p,s=\pm 1,\\ \varepsilon\in\mathcal{E}(p),\mathbf{m^{z}}\end{subarray}}\mathcal{L}^{\text{TC}^{\text{Z-loop}}}_{p,s,\varepsilon,\mathbf{m^{z}}},$
@@
Figure 5: Time evolution of flag and defect densities, together with the logical fidelity, for the Toric code on a $48\times 48$ honeycomb lattice for four representative points in the $\gamma_{x}$ - $\gamma_{z}$ phase diagram (see Fig. 1(a)). The probabilities of logical $X$ errors ($\equiv 1-F_{|\widetilde{0}\rangle_{L}}$) and logical $Z$ errors ($\equiv 1-F_{|\widetilde{+}\rangle_{L}}$) are computed by decoding the state at time $t$ using MWPM decoder (see Eqs. (15),(18)). The noise rate is set to $\eta=1$, with each data point representing an average over $500$ Monte Carlo runs. (a) Absorbing phase $(\gamma_{x}=4,\gamma_{z}=8)$: where flags rapidly reach their absorbing state $n^{f}_{X}=n^{f}_{Z}=1$, and all qubits become maximally mixed. The logical error rates $1-F_{|\widetilde{0}\rangle_{L}}$ and $1-F_{|\widetilde{+}\rangle_{L}}$ both saturate to their maximal value. (b) Partially X-active phase $(\gamma_{x}=15,\gamma_{z}=8)$: The density of $X$ flags and $B$ defects saturates at a small value, whereas $Z$ flags and $A$ defects proliferate, leading to small (large) values of logical $X(Z)$ error rate. (c) The analogous partially Z-active phase $(\gamma_{x}=4,\gamma_{z}=20)$ (d) Active phase $(\gamma_{x}=15,\gamma_{z}=20)$: here, both $X$ and $Z$ flag densities saturate to small values. Both logical $X$ (magenta colored curve) and logical $Z$ (cyan colored curve) error rates remain close to zero, and hence the quantum information can be recovered up to times exponentially large in system size.
@@ III.3 Metrics for topological order and numerical results
The full Hilbert space of our model, consisting of both qubits and flags, is spanned by the combined eigenstates of the $A_{p},B_{v}$ stabilizers and the flag occupation-numbers. The noise and correction Lindbladians in Eq. (13), which generate the total dynamics, do not couple the diagonal and off-diagonal elements of the density matrix $\rho$ in this basis. Hence, the time evolution can be simulated as a stochastic Monte Carlo dynamics of stabilizer and flag populations [81]. We initialize the system in the Toric code ground state subspace, where $B_{v}=A_{p}=+1$ for all vertices $v$ and plaquettes $p$, with the flags set to $n^{f}_{X_{j}}=n^{f}_{Z_{j}}=0$ on all edges $j$. We then simulate the dynamics using random sequential updates of the state, where the rates of all processes are specified in the previous section. The time-evolved observables are computed by averaging over multiple independent Monte Carlo runs (see Appendix. A for additional details). In what follows, we set the noise rate of the erasure channel to $\eta=1$ and explore the steady-state phase diagram as a function of the $X$ and $Z$ correction rates $\gamma_{x}$ and $\gamma_{z}$, respectively.
@@
In the active phase, the stabilizer defects are corrected along the strings of erasures frequently enough such that the density of defects remains small, as shown in Fig. 5(d) at a representative point in the phase. Here, the densities of the vertex and plaquette stabilizer defects are defined, respectively, as
@@
The heralded nature of the noise and the correction dynamics also ensures that pairs of vertex (plaquette) defects are necessarily connected by strings of $X$ ($Z$) type flags. Hence, a low density of flags necessarily corresponds to confinement of the stabilizer defects. We further illustrate this by evaluating the probability of a logical error upon pairing the defects in the time-evolved state using the MWPM decoder, ignoring the configuration of the flags. The logical state (i.e., the exact initial state within the ground state Hilbert space) is recovered by this decoder if the probability of both logical $X$ errors (logical bit-flips) and logical $Z$ errors (logical phase-flips) remains low. These error probabilities can be estimated by initializing the system in two different basis states in the Toric code ground-state space, defined as
@@
where $\mathcal{Z}_{V/H}$ are the Wilson line operators shown in Fig. 2 (which also serve as logical Pauli-$Z$ operators), and the logical Pauli-$X$ operators $\mathcal{X}_{V/H}$ are equivalent line operators comprised of products of $X$ operators along closed vertical and horizontal curves on the direct lattice. The state of the qubits in a particular Monte Carlo trajectory $r$ initialized in $|\psi_{t=0}\rangle$ can be written as
@@
[…]
@@
where $N_{r}$ is the number of independent Monte Carlo realizations. We numerically determine the state of logical information by pairing the stabilizer defects in the time-evolved state using the MWPM decoder implemented in the PyMatching-2 package [106]. The probabilities of logical $X$ errors ($\equiv 1-F_{|\widetilde{0}\rangle_{L}}$) and logical $Z$ errors ($\equiv 1-F_{|\widetilde{+}\rangle_{L}}$) are the computed in terms of the fidelity of this decoded state and two distinct initial states defined in Eq. (15). In Fig. 5(d), we see that the probability of both types of logical flips is essentially zero in the active state. The small values of logical $X$ ($Z$) error rates equivalently implies the protection of initially encoded logical $Z$ ($X$) information. This provides explicit evidence that quantum information is protected by the dynamical correction protocol for long times.
In contrast, the absorbing state is the unique dead state for the flags: when the flag density is $1$, none of our correction operations can act on any sites, and the dynamics consequently no longer acts on the flags. Once the flags enter the absorbing state, they can not leave this configuration via any of the dynamical processes. In this state, since all the correction jump operators vanish identically, the qubits evolve under the depolarizing channel and the dynamics leads to a maximally mixed steady state. This is evident in Fig. 5(a), which shows that the stabilizer defect density in this phase saturates to $1/2$. All encoded information is consequently lost, as is evident from the fact that the probability of logical flips saturates to $1-\frac{1}{2^{2}}=0.75$ since the decoder returns all possible outcomes with equal probability.
In the partially $X$-active ($Z$-active) phase, only $Z$ (only $X$) flags have proliferated. Thus, our correction protocol no longer corrects $Z$ ($X$) errors, but remains effective on $X$ ($Z$) errors. One manifestation of this is that only logical $Z$ (logical $X$) information is protected in the steady-state, as shown in Fig. 5(b) and (c). Each of these phases thus represents a classical memory, in which the information encoded in only one type of logical operator is dynamically protected. These phases represent intrinsically mixed topological phases which cannot be realized in the ground-states of 2d local gapped Hamiltonians [41, 57] since they cannot be two-way channel connected to the trivial state on the 2-torus (even though they are two-way channel connected on the infinite plane).